OpenStudy (anonymous):

Another coordinate with respect to the basis question.. ? Suppose that a rectangular x'y' coordinate system is obtained by rotating a rectangular xy coordinate system about the origin through the angle theta = 3pi/4. Find the rotations equations that express x'y'-coordinates in terms of xy coordinates and use those equations to find the x'y' coordinates of the point whose xy coordinates are (-2,6).

OpenStudy (anonymous):

I don't understand this now that it integrates theta and angles into the questions.. Can someone explain

OpenStudy (anonymous):

U want rotation matrix?

OpenStudy (anonymous):

Is that what I'm looking for? I thought it was a rotation equation

OpenStudy (anonymous):

Well it is in a way, just in matrix form..

OpenStudy (anonymous):

If u use ((cos theta, -sin theta)(sin theta, cos theta)) is a rotation through theta radians about the origin.

OpenStudy (anonymous):

I'm kind of confused b/c we're given the coordinates is it like before where we'd need vectors and a basis?

OpenStudy (anonymous):

like B={v_1, v_2,v_3} and W or something like that?

OpenStudy (anonymous):

Nothing stopping u from treating a point as a vector here. Your "basis" is the xy frame. If u apply a matrix ((a,b)(c,d)) to (x,y) u get (ax+by,cx+dy) Ç= x(ac) + y(bd)

OpenStudy (anonymous):

ignore the Ç,sorry...

OpenStudy (anonymous):

Rotating a vector and rotating the plane are "equivalent" in some sense.

OpenStudy (phi):

\[x'= xcos \theta-ysin \theta\] \[y'= xsin \theta+ycos \theta\]

OpenStudy (anonymous):

Yes, that's the matrix applied...

OpenStudy (anonymous):

blah still kind of lost, where'd you guys get these equations from

OpenStudy (anonymous):

ok all possible pairs of input can however not exist

OpenStudy (anonymous):

are those used for rotation around a vector

OpenStudy (anonymous):

No, that's a different thing....

OpenStudy (anonymous):

U understand that the equations phi posted are the ones u get from the rotation matrix?

OpenStudy (anonymous):

Yes I do those are rotation equations for the plane?

OpenStudy (anonymous):

If I rotate the plane then a vector will get carried along with it....

OpenStudy (phi):

http://en.wikipedia.org/wiki/Rotation_matrix Lots of boring stuff here!

OpenStudy (anonymous):

Or a point if u prefer (I hate that they do double duty but...).

OpenStudy (anonymous):

Oh yeah that make sense about the plane and vector thing but so if I put in the theta into the equation I get the equation right but what do I do with the coordinates

OpenStudy (anonymous):

Just put your x,y into phi's equations plus your angle and u get x',y'

OpenStudy (phi):

The coordinates (x,y) are used in the rotation eq to find (x',y')

OpenStudy (phi):

What estudier said.

OpenStudy (anonymous):

okay what do those values of coordinates represent again

OpenStudy (anonymous):

Which values?

OpenStudy (anonymous):

the x' and y' is that the new coordinates after the system is rotated?

OpenStudy (anonymous):

Yup

OpenStudy (anonymous):

Oh okay I kind of follow where you guys are going at. Thanks a lot of all your help!! Very much appreciated

OpenStudy (anonymous):

ur welcome.